OpenStudy (anonymous):

Solve the following proportion: 3 over x equals 75 over 100

OpenStudy (anonymous):

the answer 4

OpenStudy (anonymous):

3/x=75/100 now solve for x

OpenStudy (hoblos):

3/x = 75/100 cross multiply 300 = 75x x=4

OpenStudy (anonymous):

thanks you guys

OpenStudy (anonymous):

hoblos imma need your help so can you help me

OpenStudy (hoblos):

yes of course i will and im glad to XD just post the question ;)